# COMET: Co-Optimization of a CNN Model using Efficient-Hardware OBC Techniques

### v2

@@ V-A Performance Metric Definitions
TABLE III: Performance Comparison of the State-of-the-Art CNN Accelerators Based on LeNet-5 Model on FPGA
@@
Bitwidths 12/12 8/8 8/8 12/8 - 8/8 8/8 8/4 8/4 8/8 8/4 8/4
Model LeNet-5 Modified LeNet-5
Platform ZYBO Z7 ZCU106 XCVU9P Zynq-7020 Zynq-MPSoC Alpha Data 9H7 ZCU106
Freq (MHz) 100 150 300 100 100 250 100 100 95 100 100 95
LUTs 17420 514000 57657 18421 53292 16406 14724 23019 16310 14741 23071
$F$s 10195 28311 17472 1177 1044 1043 1170 1041 1041
DSPs 72 512 123 97 2614 0 0 0 0 0 0
BRAM 80 1,024 102 13.5 341 0 0 0 0 0 0
Power (W) 21.2 1.598 0.64 0.835 0.824 0.976 0.832 0.826 0.949
$\mathcal{T}_{\textnormal{MAC}}$ (GOP/s) 19.20 783.1 1490.5 0.141 3.02 110.8 0.2 0.2 0.38 0.2 0.2 0.38
ENS 21024 299918 38862 16107 320470 4102 3681 5755 4078 3685 5768
EPS 0.027 11.333 0.212 4.175 4.120 2.568 4.160 4.130 2.497
AEP 0.913 0.0166 0.0363 0.0018 0.001 0.488 0.543 0.695 0.490 0.543 0.694
  • ENS = LUT/4 + DSP $\times 102.4$ + BRAM $\times 116.2$ [36]. Assumptions: 1 DSP = $128\times 0.8=102.4$ slices, 1 BRAM (18k) = $166\times 0.7=116.2$ slices, 1 slice = 4 LUTs. EPS = Power / $\mathcal{T}_{\textnormal{MAC}}$ (W per GOP/s). AEP = $\mathcal{T}_{\textnormal{MAC}}/(f_{clk}\times\textnormal{ENS}/1000)$, reported in ops/cycle/kENS (kilo-ENS).
V-B Comparison with State-of-the-Art CNNs
@@
  • LUT usage ranges from 14,724 to 23,071 and FFs from 1,041 to 1,177, while DSPs and BRAMs are entirely eliminated. This reduction in dedicated resources brings the ENS down to just 3,681–5,768 slices, far below DSP-centric designs.
  • GOP/s varies between 0.2–0.38. Some designs operate at a slightly lower frequency of 95 MHz due to timing closure constraints on the target FPGA, yet they maintain competitive throughput relative to their resource footprint.
  • EPS lies between 2.497–4.175 W/GOP/s, and AEP ranges from 0.488–0.695 ops/cycle/kENS. These values highlight highly efficient per-slice computation as well as favorable energy utilization compared to prior art.
  • In contrast to DSP- and BRAM-intensive approaches, the proposed designs show that high normalized efficiency can be achieved entirely with LUT-based computation, without the need for higher operating frequencies or large dedicated arithmetic resources.
The results clearly demonstrate that the proposed approach offers highly efficient resource utilization for low-precision CNN on resource-constrained platforms.
Table III highlights the contrast between prior DSP- and BRAM-intensive designs and the proposed DSP and BRAM-free approach. Conventional designs achieve high absolute throughput but require large FPGA resources:
  • [16] achieves the highest reported absolute throughput of 1,490.53 GOP/s using 512 DSPs, 1,024 BRAMs, 514,000 LUTs, and unspecified FFs at 300 MHz. However, the ENS is $\approx$299,918 slices, and the AEP is only 0.0166 ops/cycle/kENS, showing that this design prioritizes peak throughput at the expense of per-resource efficiency.
  • [15] reports 19.20 GOP/s using 72 DSPs, 80 BRAMs, 17,420 LUTs, and 10,195 FFs at 100 MHz. With an ENS of $\approx$21,024 and an AEP of 0.913 ops/cycle/kENS, this design exhibits better normalized efficiency than [16], but involve DSPs and BRAMs remain underutilized for low-precision operations.
  • [38] consumes 123 DSPs, 102 BRAMs, 57,657 LUTs, and 28,311 FFs for just 0.141 GOP/s at 100 MHz. The corresponding ENS is $\approx$38,862, EPS reaches 11.333 W/GOP/s, and AEP is only 0.0363 ops/cycle/kENS—clearly illustrating high resource and power overhead with limited throughput.
  • Other reported works [37, 39] achieve moderate throughput values but still depend heavily on DSPs and BRAMs, leading to low per-slice efficiency (AEP << 0.002) despite optimizations in arithmetic precision or resource mapping.
  • Although [15] reports a higher AEP than our proposed variants, its evaluation does not include power consumption results, preventing a complete assessment of overall efficiency. Moreover, the design requires substantially more flip-flops (FFs), which directly contribute to higher dynamic power and area usage, thereby reducing its practical resource efficiency compared with our implementations.
Overall, the proposed CNN accelerator achieves a favorable trade-off between throughput, FPGA resources (LUTs, FFs, DSPs, BRAMs), and energy efficiency (EPS, AEP). The Hybrid and Split variants along with Scheme A and B provide flexibility to tune enhance efficiency of the CNN inference. This confirms that the proposed designs are well-suited for low-power, resource-constrained edge AI applications.
VI Conclusion
We presented an efficient CNN implementations based on the proposed hardware OBC techniques. By formulating CNN inference in OBC form for both inputs and weights, bit-width asymmetry has been effectively exploited and its influence on hardware efficiency analyzed. Furthermore, OBC-based LUT implementations—parallel, shared, split, and hybrid—have been developed and evaluated to highlight their trade-offs in performance, resource usage, and energy efficiency. These optimizations have been integrated into an OBC-GEMM core using the im2col transformation, enabling streamlined and resource-aware CNN execution. These methods have been encapsulated within the COMET framework, providing a unified platform for co-optimized CNN acceleration. Experimental results on a fixed-point modified LeNet-5 have demonstrated substantial reductions in resource utilization compared to state-of-the-art CNNs based on LeNet-5 [16]. By leveraging FPGA LUT fabric and eliminating reliance on DSPs and BRAMs in the compute datapath, the proposed approach offers a scalable solution for lightweight CNN deployment on edge devices. Overall, this work establishes OBC-based computation as a promising and extensible paradigm for co-optimized CNN acceleration across diverse edge AI applications.

### v3

@@ VI-A Performance Metric Definitions
TABLE III: Performance Comparison of Original/Modified LeNet-5 and All-CNN-C Models with State-of-the-Art CNNs on FPGA
LeNet-5 Design BW Platform Freq (MHz) LUTs $F$s DSPs BRAM Power (W) $\mathbf{\mathcal{T}_{\textnormal{MAC}}}$ (GOP/s) ENS EPS AEP
TCS-II [8] 12/12 ZYBO Z7 100 17420 10195 72 80 19.20 21024.00 9.13
Access [35] 8/8 ZCU106 150 21.2 1.04 20.33
TVLSI [34] 8/8 XCVU9P 300 514000 512 1024 1490.50 299918.80 16.57
Elect [5] 12/8 Zynq-7020 100 57657 28311 123 102 1.598 0.14 38861.85 11.33 0.04
A. Sci. [22] - Zynq-MPSoC 100 18421 17472 97 13.5 3.22 0.08 16106.85 40.76 0.05
Elect. [10] 8/8 Alpha Data 9H7 250 53292 2614 341 110.80 320469.80 1.38
@@
[…]
All-CNN-C CDASI [25] 16/8 Zynq-XC7Z007S 80 12610 12501 54 44 $\sim$ 2 10.62 13794.90 0.19 9.62
MAC Baseline 8/8 XCZU21DR 95 59648 1228 27 64 1.662 0.19 25113.60 8.75 0.08
MAC Baseline 8/4 95 59086 1081 27 32 1.598 0.38 21254.70 4.21 0.19
[…]
  • BW: Bitwidth format (input act./weight), $^{\clubsuit}:$ Modified LeNet-5, $^{\spadesuit}:$ with BRAM, ENS = LUT/4 + DSP $\times 102.4$ + BRAM $\times 116.2$ [23]. Assumptions: 1 DSP = $128\times 0.8=102.4$ slices, 1 BRAM (18k) = $166\times 0.7=116.2$ slices, 1 slice = 4 LUTs. EPS = Power / $\mathcal{T}_{\textnormal{MAC}}$ (W per GOP/s). AEP = $1000\times\mathcal{T}_{\textnormal{MAC}}/(f_{clk}\times\textnormal{ENS}/1000)$, reported in ops/cycle/kiloENS.
VI-B Evaluation of the Models and their Comparison
[…]
The original LeNet-5 maintains a consistent 90 MHz across all variants. For the modified LeNet-5 on the ZCU106, Scheme A variants operate at 100 MHz, while Scheme B variants operate at 95 MHz. Both the original and modified LeNet-5 designs eliminate DSPs and BRAMs, relying solely on slice LUTs. The modified LeNet-5 achieves a substantial reduction in LUT utilisation and power consumption (0.824–0.976 W vs. 1.41–1.52 W), though at the cost of a modest increase in FF count. Throughput remains comparable across both variants, while Scheme B provides higher computational speed than Scheme A, at the cost of higher LUT usage and greater bandwidth demand to sustain the increased processing rate. The modified LeNet-5 demonstrates improved energy efficiency with lower EPS values (Hybrid and Split: 2.57–4.18 vs. 4.17–8.00 W/GOP/s), while it also achieves higher AEP (Hybrid and Split: 0.49–0.70 vs. 0.15–0.24 ops/cycle/kENS), indicating better throughput per cycle relative to its resource footprint. It is therefore evident that the modified LeNet-5 offers a more favourable balance between resource utilisation and energy efficiency, making it better suited for resource-constrained deployment when implemented using slice LUTs only. Table III also compares the original and modified LeNet-5 designs with state-of-the-art implementations. Prior work such as [34] achieves very high throughput of 1490.5 GOP/s at 300 MHz by relying heavily on DSPs and BRAMs, while smaller-scale implementations such as [5] provide only modest throughput at considerable resource and energy cost. In contrast, both the original and modified LeNet-5 designs deliver competitive EPS against other MAC-based implementations. While a direct AEP comparison with MAC-based implementations is inherently less straightforward due to the bit-serial nature of OBC schemes, the modified LeNet-5 nevertheless achieves superior AEP over several smaller-scale implementations — demonstrating a favourable balance between computational efficiency and resource cost.
In contrast, All-CNN-C is evaluated under two configurations. For the BRAM-free implementation, the network parameters must be mapped entirely onto slice resources, imposing a significantly higher LUT requirement; therefore, the LUT-richer XCZU21DR is selected, with a maximum frequency of 50 MHz. For the BRAM-enabled implementation, main storage is absorbed by on-chip block memory, significantly reducing LUT pressure; as a result, the design can be implemented on the resource-constrained ZCU106, operating at  100 MHz. This comparison is used to evaluate whether a slice-LUT-only or BRAM-assisted implementation is more suitable for the deeper All-CNN-C network. The results in Table III show that, compared with the slice-LUT-only XCZU21DR configuration, the BRAM-enabled ZCU106 configuration operates at a higher clock frequency with lower power consumption, yielding significantly improved EPS (Hybrid and Split: 3.09–6.86 vs. 25.79–43.50 W/GOP/s) and AEP (Hybrid and Split: 0.089–0.192 vs. 0.024–0.040 ops/cycle/kENS). This suggests that, as network complexity increases, BRAM utilisation helps alleviate LUT pressure and improve overall energy efficiency [30]. This contrasts with the LeNet-5 findings, where the slice-LUT-only approach proved sufficient and effective. For All-CNN-C [29], only a limited number of FPGA implementations are available in the literature. The only directly comparable reference is [25], which implements the same network but reports an inaccurate power figure of $\sim$2 W from the Xilinx Power Estimator and thus the derived EPS value is indicative rather than definitive. MAC baselines of All-CNN-C are therefore also realized, employing DSPs and BRAMs, consuming 1.60–1.66 W compared to 1.23–1.37 W for the proposed BRAM-enabled ZCU106 variants. The proposed variants operate entirely without DSPs, yet achieve comparable EPS (Hybrid and Split: 3.09–6.86 vs. 4.21–8.75 W/GOP/s) and AEP (0.089–0.192 vs. 0.080–0.188 ops/cycle/kENS) to the MAC baselines, demonstrating that DSP-free LUT-based inference remains competitive in energy efficiency for deeper networks, despite the inherent routing complexity and clock frequency limitations associated without BRAM exploitation.
VII Conclusion
We have presented COMET, a co-optimization framework for DSP-free CNN inference on FPGAs using OBC techniques. Inference has been formulated with OBC representations applied separately to inputs (Scheme A) and weights (Scheme B), exploiting bit-width asymmetry, while the SA operation incorporates the offset term with pre-scaled bias. OBC-based LUT implementations, including parallel, shared, split, and hybrid, have been evaluated to highlight trade-offs in performance, resource usage, and energy efficiency, and integrated into an OBC-GEMM core via the im2col transformation for resource-aware CNN execution. Both the original and modified LeNet-5 have demonstrated the viability of slice-LUT-only CNN inference, with the modified design achieving a more favourable balance between resource utilisation and energy efficiency. All-CNN-C has further validated scalability for deeper networks, where BRAM utilisation has been shown to alleviate LUT pressure and improve energy efficiency as network complexity increases. These results collectively demonstrate that DSP-free FPGA deployment is achievable across CNN models of varying complexity, establishing COMET as a scalable and practical solution for LUT-centric CNN inference across edge AI applications.